# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 6.1. Elementary flops
As $\mathcal{X}_{0}$ is generic as in Definition 3.19, and so does not contain any internal $(-2)$-curve, it follows from the proof of [86, Lemma 3.2] that the connected components of the exceptional locus of $b$ are either as in i) or ii). Hence, to prove Lemma 6.1, it is enough to show that the exceptional locus of $b$ is in fact connected. Assume that $C$ and $C^{\prime}$ are two curves as in i) or ii) with proportional classes. Curves as in i) have negative intersection with exactly one irreducible component of $\mathcal{X}_{0}$, whereas curves as in ii) have negative intersection with exactly two irreducible components of $\mathcal{X}_{0}$. Hence, either both $C$ and $C^{\prime}$ are as in i), or both $C$ and $C^{\prime}$ are as in ii). In the latter case, we have $C=C^{\prime}=V_{i}\cap V_{j}$, where $V_{i}$ and $V_{j}$ are the two irreducible components of $\mathcal{X}_{0}$ with $C\cdot V_{i}<0$ and $C\cdot V_{j}<0$. In the former case, both $C$ and $C^{\prime}$ are internal exceptional curves in the same irreducible component $V_{i}$ of $\mathcal{X}_{0}$, the one such that $C\cdot V_{i}<0$, and intersect the same irreducible component $V_{i}\cap V_{j}$ of $\partial V_{i}$: $V_{j}$ is the unique irreducible component of $\mathcal{X}_{0}$ such that $C.V_{i}>0$. Hence, the line bundle $\mathcal{O}_{V_{i}}(C-C^{\prime})$ is trivial in restriction to every irreducible component of $\partial V_{i}$, and so, by the description of $\mathrm{Pic}(\mathcal{X}_{0})$ given by Lemma 3.10, defines a line bundle on $\mathcal{X}_{0}$. By Lemma 3.10, we also have $H^{2}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and so this line bundle formally deforms into a line bundle $L$ on $\mathcal{X}$. If $C\neq C^{\prime}$, then $L\cdot C=-1$, whereas $L\cdot C^{\prime}=1$, and so, as the classes $C$ and $C^{\prime}$ are assumed to be proportional, $C=-C^{\prime}$, in contradiction with the strict convexity of the cone of curves $NE(\mathcal{X}/\overline{\mathcal{X}})$. ∎
Following [7, Definition 7.15], we call M1 flop a small $\mathbb{Q}$-factorial modification as in Lemma 6.1 i), that is the flop of an internal exceptional curve, and M2 flop a small $\mathbb{Q}$-factorial modification as in Lemma 6.1 ii), that is the flop of an exceptional curve contained in the double locus. In restriction to the central fiber $\mathcal{X}_{0}$, M1 and M2 flops act like the M1 and M2 modifications of open Kulikov surfaces introduced in §3.3. Note that M1 and M2 flops are often referred to as Type 1 and Type 2 modifications in the literature [87, pg. 12].
@@ 6.2. Structure constants under flops
In this section, we describe the behavior of structure constants $N_{pqr}^{\beta}(\mathcal{X})$ under M1-M2 flops. Parallel results are obtained in [44] in the context of Type III degenerations of K3 surfaces.
Since M1 and M2 flops are isomorphisms in codimension one, the Picard group of an open Kulikov degeneration remains invariant under such flops. Thus we have a natural isomorphism
$h\colon N_{1}(\mathcal{X}/\overline{\mathcal{X}})\longrightarrow N_{1}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$
where $N_{1}(\mathcal{X}/\overline{\mathcal{X}})$ denotes the group of numerical equivalence classes of one cycles in $\mathcal{X}$, whose irreducible components are curves contracted to a point under the map $f\colon\mathcal{X}\to\overline{\mathcal{X}}$. Here, by the definition of M1 and M2 flops, the image of the flopping curve class $[C]\in N_{1}(\mathcal{X}/\overline{\mathcal{X}})$ under the map in Equation (6.2) is $-[C^{\prime}]\in N_{1}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$, where $C^{\prime}$ is the flopped curve. Moreover, there is a natural identification between the integral points of the tropicalization $C(P)$ of $\mathcal{X}\to\Delta$ and the integral points of the tropicalization $C(P^{\prime})$ of $\mathcal{X}^{\prime}\to\Delta$.
@@
Let $\mathcal{X}\rightarrow\Delta$ and $\mathcal{X}^{\prime}\to\Delta$ be two quasi-projective open Kulikov degenerations differing by an M1 flop. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}\cong C(P^{\prime})_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$, we have the equality $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{h(\beta)}(\mathcal{X}^{\prime})$, where $h$ is the isomorphism given in Equation (6.2).
@@
It follows from the properness of moduli spaces given by Theorem 5.2 that the results proved in [40, §8] on punctured Gromov–Witten invariants of Type III degenerations of K3 surfaces also hold by the same arguments for the punctured Gromov–Witten $N_{pqr}^{\beta}(\mathcal{X})$ of quasi-projective open Kulikov degenerations. By [40, Theorem 8.15], the punctured Gromov–Witten invariants $N_{pqr}^{\beta}(\mathcal{X})$ can be decomposed as a product of punctured Gromov–Witten invariants of the irreducible components $(X^{i},\partial X^{i})$ of $\mathcal{X}_{0}$ and of toric surfaces (appearing as irreducible components of a large enough base change of $\mathcal{X}\rightarrow\Delta$). On the other hand, the M1 flop $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ induces an isomorphism between $\mathcal{X}_{0}$ and $\mathcal{X}_{0}^{\prime}$, except in restriction to the irreducible component $X^{v}$ containing the flopping curve $C$, and to the only other irreducible components $X^{v^{\prime}}$ intersecting $C$. Hence, Theorem 6.2 reduces to a local statement comparing punctured Gromov–Witten invariants of $X^{v}\cup X^{v^{\prime}}$ under the blow-down of $C$ in $X_{v}$, and the blow-up of the point $C\cap X^{v^{\prime}}$ in $X^{v^{\prime}}$. Finally, this statement follows from [40, Theorem 8.16] describing the change of the punctured Gromov–Witten invariants of a log Calabi–Yau surfaces under internal blow-up. We refer to [44] for more details. ∎
@@
An M1 flop induces an isomorphism at the level of polyhedral complexes $(P,\mathscr{P})$. At the level of the integral affine structure with singularities on $P$, a focus-focus singularity moves along its monodromy invariant from the vertex $v$, corresponding to the component $X^{v}$ containing the flopping curve $C$, to the vertex $v^{\prime}$, corresponding to the component $X^{v^{\prime}}$ containing the flopped curve $C^{\prime}$.
Let $\mathcal{X}^{\prime}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration obtained from a quasi-projective open Kulikov degeneration $\mathcal{X}\rightarrow\Delta$ by an M2 flop along a curve $C$ in $\mathcal{X}$. The curve $C$ corresponds to an edge $e_{C}$ of $\mathscr{P}$, which is a diagonal in a quadrilateral $Q$. The polyhedral decomposition $\mathscr{P}^{\prime}$ of $P$ defined by $\mathcal{X}^{\prime}$ is obtained from $\mathscr{P}$ by flip of $e_{C}$ in $Q$, that is, by removing the diagonal $e_{C}$ of $Q$ and introducing the opposite diagonal $e_{C^{\prime}}$ of $Q$. Let $c_{Q}\in C(\mathscr{P})_{\mathbb{Z}}$ be the integral primitive generator of the ray in $C(\mathscr{P})$ passing through the intersection point of $e_{C}$ and $e_{C^{\prime}}$ in $P\subset C(P)$. Geometrically, $c_{Q}$ is the divisorial valuation defined by the exceptional divisor $E$ of the blow-up $\mathcal{Z}$ of the ordinary double point obtained by contracting $C$. Denote by $\widetilde{C(\mathscr{P})}$ the dual intersection complex of $\mathcal{Z}$, that is, the cone complex obtained from $C(\mathscr{P})$ by adding the cone over $e_{C^{\prime}}$. Let
$\varphi_{E}:C(P)\longrightarrow\mathbb{R}$ (6.1)
be the unique piecewise-linear function, linear on each cone of $\widetilde{C(\mathscr{P})}$, such that $\varphi_{E}(p)=0$ for every $p\in P_{\mathbb{Z}}$ and $\varphi_{E}(c_{Q})=1$.
@@
Let $\mathcal{X}^{\prime}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration obtained from a quasi-projective open Kulikov degeneration $\mathcal{X}\rightarrow\Delta$ by an M2 flop along a curve $C$ in $\mathcal{X}$. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}\cong C(P^{\prime})_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$, we have the equality
$N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{h(\beta)+m_{pqr}[C]}(\mathcal{X}^{\prime})\,,$
@@
$m_{pqr}=\varphi_{E}(p)+\varphi_{E}(q)-\varphi_{E}(r)\,,$
where $\varphi_{E}$ is the PL function as in Equation (6.1).
@@
The M2 flop is defined by a diagram
@@
As $\mathcal{Z}$ is a log modification of $\mathcal{X}$, induced by the subdivision $\widetilde{C(\mathscr{P})}$ of $C(\mathscr{P})$, we have by [56, §10]
$N_{pqr}^{\beta}(\mathcal{X})=\sum_{\begin{subarray}{c}\widetilde{\beta}\in NE(\mathcal{Z}/\overline{\mathcal{X}})\\ p_{\star}\widetilde{\beta}=\beta\end{subarray}}N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})\,.$
In fact, arguing as in the proof of [40, Lemma 8.7], there exists a unique $\widetilde{\beta}$ such that $N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})\neq 0$, and so we have $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})$. Similarly, we have $N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})=N_{pqr}^{p^{\prime}_{\star}\widetilde{\beta}}(\mathcal{X}^{\prime})$. Hence, in order to prove Theorem 6.4, it is enough to show that $p^{\prime}_{\star}\widetilde{\beta}=h(\beta)+m_{pqr}C$. This formula holds since, by geometry of the flop, we have $p^{\prime}_{\star}\widetilde{\beta}=h(\beta)+(E\cdot\widetilde{\beta})C$, and we have $E\cdot\widetilde{\beta}=\varphi_{E}(p)+\varphi_{E}(q)-\varphi_{E}(r)$ by Lemma 6.7. ∎
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. As in [45, Definition 1.8], let $\varphi$ be the $N_{1}(\mathcal{X}/\overline{\mathcal{X}})_{\mathbb{R}}$-valued multivalued PL function on $C(P)$ with kink $[C]\in NE(\mathcal{X}/\overline{\mathcal{X}})$ across the codimension one face of $C(\mathscr{P})$ corresponding to the irreducible component $[C]$ of the double locus of $\mathcal{X}_{0}$. By [45, Construction 1.14], $\varphi$ determines a cone complex $(C(P)_{\varphi},C(\mathscr{P})_{\varphi})$, and a map of cone complexes $\pi:(C(P)_{\varphi},C(\mathscr{P})_{\varphi})\rightarrow(C(P),C(\mathscr{P}))$ such that $\pi:C(P)_{\varphi}\rightarrow C(P)$ is a $N_{1}(\mathcal{X}/\overline{\mathcal{X}})_{\mathbb{R}}$-torsor, and $\varphi$ can naturally be viewed as a section of $\pi$.
Let $\mathcal{X}\rightarrow\Delta$ and $\mathcal{X}^{\prime}\rightarrow\Delta$ be two quasi-projective open Kulikov degenerations with $\overline{\mathcal{X}}=\overline{\mathcal{X}}^{\prime}$, and $\varphi$ and $\varphi^{\prime}$ the corresponding multivalued PL functions. It follows from [50, §2.1] that we have a natural homeomorphism $C(P)_{\varphi}\simeq C(P^{\prime})_{\varphi^{\prime}}$, preserving the integral points, but not the cone decompositions $C(\mathscr{P})_{\varphi}$ and $C(\mathscr{P}^{\prime})_{\varphi^{\prime}}$ in general. In particular, the difference $\varphi^{\prime}-\varphi$ can be viewed as an ordinary $N_{1}(\mathcal{X}/\overline{\mathcal{X}})_{\mathbb{R}}$-valued function on $C(P)_{\mathbb{Z}}\simeq C(P^{\prime})_{\mathbb{Z}}$.
[…]
Figure 6.1. Change of the polyhedral decomposition $\mathscr{P}$ under a M2 flop on the left. The function $\varphi_{E}$ on the right.
@@
Let $\mathcal{X}^{\prime}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration obtained from a quasi-projective open Kulikov degeneration $\mathcal{X}\rightarrow\Delta$ by an M2 flop along a curve $C$ in $\mathcal{X}$. Then, for every $p\in C(P)_{\mathbb{Z}}\simeq C(P^{\prime})_{\mathbb{Z}}$, we have
$\varphi^{\prime}(p)-\varphi(p)=\varphi_{E}(p)[C]\,,$
where $\varphi_{E}$ is the PL function as in Equation (6.1).
@@
Recall that $\mathscr{P}^{\prime}$ is obtained from $\mathscr{P}$ by flipping a diagonal $e_{C}$ of a quadrilateral $Q$ into the opposite diagonal $e_{C^{\prime}}$ – see Figure 6.1. Away from $Q$, the decompositions $\mathscr{P}$ and $\mathscr{P}^{\prime}$ coincide, we have $\varphi_{E}=0$, and so we clearly have $\varphi^{\prime}=\varphi+\varphi_{E}[C]$. Hence, it is enough to show that $\varphi+\varphi_{E}[C]$ has the same kinks as $\varphi^{\prime}$ in restriction to a neighborhood of $Q$. Identifying $Q$ with the standard square of integral size one, the restriction of $\varphi_{E}$ to $Q$ is given as in Figure 6.1, and so has kinks $1$ across the sides of $Q$ and kinks $-1$ across the 4 connected components of $(e_{C}\cup e_{C^{\prime}})\setminus(e_{C}\cap e_{C^{\prime}})$. In particular, given a side $s$ of $Q$ corresponding to an irreducible component $C_{s}$ of the irreducible component of the double locus of $\mathcal{X}_{0}$, the kink of $\varphi+\varphi_{E}[C]$ across $s$ is $[C_{s}]+[C]=[C_{s}]-[C^{\prime}]$, which is exactly the class of the strict transform of $C_{s}$ under the M2 flop, and so the kink of $\varphi^{\prime}$ across $s$. Finally, the kink of $\varphi+\varphi_{E}[C]$ across $e_{C}$ (resp. $e_{C^{\prime}}$) is $[C]-[C]=0$ (resp . $0-[C]=[C^{\prime}]$), which is exactly the corresponding kink of $\varphi^{\prime}$. ∎
@@
Let $\mathcal{X}\rightarrow\Delta$ and $\mathcal{X}^{\prime}\rightarrow\Delta$ be two quasi-projective open Kulikov degenerations with $\overline{\mathcal{X}}=\overline{\mathcal{X}}^{\prime}$. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}\cong C(P^{\prime})_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$, we have the equality
@@
$n_{pqr}=(\varphi^{\prime}(p)-\varphi(p))+(\varphi^{\prime}(q)-\varphi(q))-(\varphi^{\prime}(r)-\varphi(r))\in N_{1}(\mathcal{X}/\overline{\mathcal{X}})\simeq N_{1}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})\,.$

### v3

@@ 6.1. Flops of relative Picard rank one
It remains to treat the case of two disjoint interior exceptional curves $C_{1}$ and $C_{2}$. As the irreducible components of $\mathcal{X}_{0}$ are normal, there exists $v$ and $w\neq v$ such that $C_{1}\subset X^{v}$, $C_{1}\cdot X^{v}=-1$, and $C_{1}\cdot X^{w}=1$. Moreover, $X^{v}$ and $X^{w}$ are exactly the irreducible components of $\mathcal{X}_{0}$ that have non-zero intersection with $C_{1}$. If $C_{2}$ is positively proportional to $C_{1}$, then we also have $C_{2}\cdot X^{v}=-1$, hence $C_{2}\subset X^{v}$, and $C_{2}\cdot X^{w}=1$. Since the double intersection $X^{v}\cap X^{w}$ is connected by assumption, the line bundle $\mathcal{O}_{X^{v}}(C_{1}-C_{2})$ is trivial in restriction to every irreducible component of $\partial X^{v}$, and so, by the description of $\mathrm{Pic}(\mathcal{X}_{0})$ given by Lemma 3.10, defines a line bundle on $\mathcal{X}_{0}$. By Lemma 3.10, we also have $H^{2}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}})=0$, and so this line bundle formally deforms into a line bundle $L$ on $\mathcal{X}$. As $C_{1}$ and $C_{2}$ are disjoint curves, we obtain that $L\cdot C_{1}=-1$ and $L\cdot C_{2}=1$. Since the classes $C_{1}$ and $C_{2}$ are assumed to be proportional, we deduce that $C_{1}=-C_{2}$, in contradiction with the strict convexity of the cone of curves $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. ∎
@@ 6.2. Structure constants under flops
In this section, we describe the behavior of structure constants $N_{pqr}^{\beta}(\mathcal{X})$ under flops. Parallel results are obtained in [47] in the context of Type III degenerations of K3 surfaces, and we refer to [47] for a more detailed exposition.
Since a flop $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ is an isomorphism in codimension one, the Picard groups of $\mathcal{X}$ and $\mathcal{X}^{\prime}$ are naturally identified. Consequently, we obtain a natural isomorphism
$h\colon N_{1}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\longrightarrow N_{1}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})\,$, (6.1)
where $N_{1}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ denotes the group of numerical equivalence classes of one cycles in $\mathcal{X}$, whose irreducible components are curves contracted to a point under the map $f\colon\mathcal{X}\to\mathcal{X}^{\mathrm{can}}$. For instance, if $C$ is a flopping curve in $\mathcal{X}$ for a multiple M1 or M2 flop, then we have $h([C])=-[C^{\prime}]$, where $C^{\prime}$ is the flopped curve in $\mathcal{X}^{\prime}$. Additionally, there is a natural identification $C(P)_{\mathbb{Z}}\simeq C(P^{\prime})_{\mathbb{Z}}$ between the integral points of the tropicalization $C(P)$ and $C(P^{\prime})$ of $\mathcal{X}\to\Delta$ and $\mathcal{X}^{\prime}\to\Delta$ respectively.
We now describe the transformation of structure constants $N_{pqr}^{\beta}(\mathcal{X})$ under multiple M1 or M2 flops.
@@
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration and $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ be a multiple M1 flop. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}\cong C(P^{\prime})_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, we have the equality $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{h(\beta)}(\mathcal{X}^{\prime})$, where $h$ is the isomorphism given in Equation (6.1).
@@
It follows from the properness of moduli spaces given by Theorem 5.2 that the results proved in [43, §8] on punctured Gromov–Witten invariants of Type III degenerations of K3 surfaces also hold by the same arguments for the punctured Gromov–Witten $N_{pqr}^{\beta}(\mathcal{X})$ of quasi-projective open Kulikov degenerations. By [43, Theorem 8.15], the punctured Gromov–Witten invariants $N_{pqr}^{\beta}(\mathcal{X})$ can be decomposed as a product of punctured Gromov–Witten invariants of the irreducible components of $\mathcal{X}_{0}$ and of toric surfaces appearing as irreducible components of a resolution of a large enough base change of $\mathcal{X}\rightarrow\Delta$. On the other hand, a multiple M1 flop only changes the irreducible components of $\mathcal{X}_{0}$ by interior blow-ups or interior blow-downs. Thus, the statement follows from [43, Theorem 8.16] describing the change of the punctured Gromov–Witten invariants of a log Calabi–Yau surfaces under interior blow-up. We refer to [47] for more details. ∎
@@
An M1 flop induces an isomorphism at the level of polyhedral complexes $(P,\mathscr{P})$. At the level of the integral affine structure with singularities on $P$, a focus-focus singularity moves along its monodromy invariant direction from the vertex $v$, corresponding to the component $X^{v}$ containing the flopping curve $C$, to the vertex $v^{\prime}$, corresponding to the component $X^{v^{\prime}}$ containing the flopped curve $C^{\prime}$.
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration and $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ be a multiple M2 flop along disjoint (1,1)(-1,-1)-double curves $C_{i}$ in $\mathcal{X}$. For every $i$, the double curve $C_{i}$ corresponds to an edge $e_{C_{i}}$ of $\mathscr{P}$, which is a diagonal in a quadrilateral $Q_{i}$. Since the double curves $C_{i}$ are disjoint, the quadrilaterals $Q_{i}$ do not overlap. The polyhedral decomposition $\mathscr{P}^{\prime}$ of $P$ defined by $\mathcal{X}^{\prime}$ is obtained from $\mathscr{P}$ by flipping $e_{C_{i}}$ in $Q_{i}$ for every $i$, that is, by removing the diagonal $e_{C_{i}}$ of $Q$ and introducing the opposite diagonal $e_{C_{i}^{\prime}}$ of $Q$. Let $c_{Q_{i}}\in C(\mathscr{P})_{\mathbb{Z}}$ be the integral primitive generator of the ray in $C(\mathscr{P})$ that passes through the intersection point of $e_{C_{i}}$ and $e_{C_{i}^{\prime}}$ in $P\subset C(P)$. Geometrically, $c_{Q_{i}}$ corresponds to the divisorial valuation defined by the exceptional divisor $E_{i}$ of the blow-up $\mathcal{Z}$ of the ordinary double point obtained by contracting $C_{i}$. Denote by $\widetilde{C(\mathscr{P})}$ the dual intersection complex of $\mathcal{Z}$, that is, the cone complex obtained from $C(\mathscr{P})$ by adding the cones over $e_{C_{i}^{\prime}}$. For every $i$, define
$\varphi_{E_{i}}:C(P)\longrightarrow\mathbb{R}$ (6.2)
as the unique piecewise linear function, linear on each cone of $\widetilde{C(\mathscr{P})}$, satisfying $\varphi_{E_{i}}(p)=0$ for every $p\in P_{\mathbb{Z}}$ and $\varphi_{E_{i}}(c_{Q_{i}})=1$ – see Figure 9.
Figure 9. Change of the polyhedral decomposition $\mathscr{P}$ under a M2 flop on the left. The function $\varphi_{E}$ on the right.
@@
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration and $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ be a multiple M2 flop along disjoint (1,1)(-1,-1)-double curves $C_{i}$ in $\mathcal{X}$. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}\cong C(P^{\prime})_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, we have the equality
$N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{h(\beta)+\sum_{i}m_{pqr,i}[C_{i}]}(\mathcal{X}^{\prime})\,$,
@@
$m_{pqr,i}=\varphi_{E_{i}}(p)+\varphi_{E_{i}}(q)-\varphi_{E_{i}}(r)\,$,
where $\varphi_{E_{i}}$ is the piecewise linear function defined in Equation (6.2).
@@
The multiple M2 flop is defined by a diagram
$\mathcal{Z}$$\mathcal{X}$$\mathcal{X}^{\prime}\,.$$\scriptstyle{p^{\prime}}$$\scriptstyle{p}$
As $\mathcal{Z}$ is a log modification of $\mathcal{X}$, induced by the subdivision $\widetilde{C(\mathscr{P})}$ of $C(\mathscr{P})$, we have by [61, §10]
$N_{pqr}^{\beta}(\mathcal{X})=\sum_{\begin{subarray}{c}\widetilde{\beta}\in NE(\mathcal{Z}/\mathcal{X}^{\mathrm{can}})\\ p_{\star}\widetilde{\beta}=\beta\end{subarray}}N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})\,.$
In fact, arguing as in the proof of [43, Lemma 8.7], there exists a unique $\widetilde{\beta}$ such that $N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})\neq 0$, and so we have $N_{pqr}^{\beta}(\mathcal{X})=N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})$. Similarly, we have $N_{pqr}^{\widetilde{\beta}}(\mathcal{Z})=N_{pqr}^{p^{\prime}_{\star}\widetilde{\beta}}(\mathcal{X}^{\prime})$. Hence, in order to prove Theorem 6.11, it suffices to show that $p^{\prime}_{\star}\widetilde{\beta}=h(\beta)+\sum_{i}m_{pqr,i}C_{i}$. This formula holds since, by geometry of the flop, we have $p^{\prime}_{\star}\widetilde{\beta}=h(\beta)+\sum_{i}(E_{i}\cdot\widetilde{\beta})C_{i}$, and we have $E_{i}\cdot\widetilde{\beta}=\varphi_{E_{i}}(p)+\varphi_{E_{i}}(q)-\varphi_{E_{i}}(r)$ by [51, Corollary 1.14]. ∎
We now aim to formulate how the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ transform under general flops. To achieve this, we first apply [48, Construction 1.14]. Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. Following [48, Construction 1.14], we define
$\mathbb{B}:=C(P)\times N_{1}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})_{\mathbb{R}}$
and equip $\mathbb{B}$ with the polyhedral decomposition
$\mathscr{P}_{\mathbb{B}}:=\{\tau\times N_{1}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})_{\mathbb{R}}\,|\,\tau\in C(\mathscr{P})\}\,.$
Additionally, there exists a natural lift of the integral affine structure defined on the complement in $C(P)$ of codimension $2$ cells of $C(\mathscr{P})$ to an integral affine structure on the complement in $\mathbb{B}$ of the codimension $2$ cells of $\mathscr{P}_{\mathbb{B}}$, see [48, Eq 1.9] : given two maximal cells $\sigma$, $\sigma^{\prime}$ of $C(\mathscr{P})$ that intersect along a codimension cell $\rho$, and a chart $f:\sigma\cup\sigma^{\prime}\rightarrow\mathbb{R}^{n}$ of the integral affine structure on $C(P)$, a corresponding chart of the integral affine structure on $\mathbb{B}$ is given by
$(\sigma\cup\sigma^{\prime})\times N_{1}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})_{\mathbb{R}}$ $\longrightarrow\mathbb{R}^{n}\times N_{1}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})_{\mathbb{R}}$ (6.3)
@@
where $\delta$ is the signed integral distance to $\rho$ that is positive on $\sigma^{\prime}$, and $C_{\rho}]\in N_{1}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ is the curve class of the double curve in $\mathcal{X}_{0}$ corresponding to the edge $\rho\cap P$ of $\mathscr{P}$. There is a natural linear projection $\pi:\mathbb{B}\rightarrow C(P)$ making $\mathbb{B}$ a $N_{1}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})_{\mathbb{R}}$-torsor over $C(P)$. Finally, we denote by
$\varphi:C(P)\longrightarrow\mathbb{B}$ (6.4)
the zero section. For the integral affine structure on $\mathbb{B}$ defined by (6.3), $\varphi$ is not a linear map but only a piecewise linear map.
A non-Archimedean interpretation of $\mathbb{B}$, $\pi$, and $\varphi$ is provided in [53, §2.1] in terms of the essential skeleton of the universal $\mathbb{G}_{m}$-torsor over the non-Archimedean generic fiber of $\mathcal{X}\rightarrow\Delta$. According to [53, §2.1], every flop $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ induces a natural piecewise linear isomorphism
$\nu_{b}:\mathbb{B}\xrightarrow{\sim}\mathbb{B}^{\prime}$
between the corresponding $N_{1}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})_{\mathbb{R}}\simeq N_{1}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})_{\mathbb{R}}$ torsors. This isomorphism does not preserve the polyhedral decompositions in general. In particular, the difference
$\varphi^{\prime}-\nu_{b}\circ\varphi$ (6.5)
can be regarded as an ordinary $N_{1}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})_{\mathbb{R}}$-valued function on $C(P)_{\mathbb{Z}}\simeq C(P^{\prime})_{\mathbb{Z}}$. The following lemmas compute this function for multiple M1 and M2 flops.
@@
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration and $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ be a multiple M1 flop. Then, we have $\varphi^{\prime}-\nu_{b}\circ\varphi=0$.
@@
Since the exceptional locus of a multiple M1 flop contains no stratum of $\mathcal{X}_{0}$, the result follows from [53, Lemma 2.10] and [53, Lemma 2.11]. ∎
Lemma 6.13.
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration and $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ be a multiple M2 flop along disjoint (1,1)(-1,-1)-double curves $C_{i}$ in $\mathcal{X}$. Then, we have
$\varphi^{\prime}-\nu_{b}\circ\varphi=\sum_{i}\varphi_{E_{i}}[C_{i}]\,$,
where $\varphi_{E_{i}}$ are the piecewise linear functions defined by Equation (6.2).
Proof.
The result follows from [53, Lemma 2.10] calculating $\varphi^{\prime}-\nu_{b}\circ\varphi$ in terms of the exceptional divisors of a diagram resolving the flop. ∎
@@
Let $\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration and $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ be a flop. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}\cong C(P^{\prime})_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, we have the equality
@@
$n_{pqr}=(\varphi^{\prime}(p)-\nu_{b}(\varphi(p)))+(\varphi^{\prime}(q)-\nu_{b}(\varphi(q)))-(\varphi^{\prime}(r)-\nu_{b}(\varphi(r)))\in N_{1}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\simeq N_{1}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})\,.$